# What’s the Weight?

### v1

@@ Abstract
A basic descriptive question in statistics often asks whether there are differences in mean outcomes between groups based on levels of a discrete covariate (e.g., racial disparities in health outcomes). However, when this categorical covariate of interest is correlated with other factors related to the outcome, direct comparisons may lead to biased estimates and invalid inferential conclusions without appropriate adjustment. Propensity score methods are broadly employed with observational data as a tool to achieve covariate balance, but how to implement them in complex surveys is less studied – in particular, when the survey weights depend on the group variable under comparison. In this work, we focus on a specific example when sample selection depends on race. We propose identification formulas to properly estimate the average controlled difference (ACD) in outcomes between Black and White individuals, with appropriate weighting for covariate imbalance across the two racial groups and generalizability. Via extensive simulation, we show that our proposed methods outperform traditional analytic approaches in terms of bias, mean squared error, and coverage when estimating the ACD for our setting of interest. We are motivated by the interplay between race and social determinants of health when estimating racial differences in telomere length using data from the National Health and Nutrition Examination Survey (NHANES). The NHANES sampling scheme depends on self-reported race. We build a propensity for race to properly adjust for other social determinants while characterizing the controlled effect of race on telomere length. We find that evidence of racial differences in telomere length between Black and White individuals attenuates after accounting for confounding by socioeconomic factors and after utilizing appropriate propensity score and survey weighting techniques. Software to implement these methods and code to reproduce our results can be found in the R package svycdiff at https://github.com/salernos/svycdiff.
Keywords Controlled Outcome Differences  $\cdot$ Propensity Scores  $\cdot$ Complex Surveys  $\cdot$ NHANES  $\cdot$ Racial Disparities
@@ 1 Introduction
There is a deep and broad literature in the United States (US) documenting racial disparities along multiple socioeconomic axes (e.g., income, education; see Needham et al., 2013; Rehkopf and Needham, 2019; Zota and VanNoy, 2021). These dimensions are, of course, not independent, making it difficult to separate the variation in outcomes associated with race from differences attributable to other sources. Self-identified race, as measured in many studies, is a complex composite of biology and lived experiences, making it impossible to disentangle the two without additional data (e.g., genetic ancestry). Propensity scores are commonly used in causal inference, but they are also applicable under weaker assumptions to study such differences among groups by balancing complex correlates across these groups (Stuart et al., 2011; Lim et al., 2018; Zeng et al., 2021). While previous work has shown that propensity scores can be used to study racial disparities (see Li and Li, 2023), a gap remains in determining the ideal practice when the sampling design is complex. Specifically, surveys are the backbone of such observational research, but, especially as response rates decline and costs rise, they almost always involve complex designs and non-response patterns. Of particular interest is when the design probability of selection into the survey sample depends on the comparison variable of interest, such as race. In more traditional causal inference terminology, these sensitive variables may be analogous to ‘treatment’ or ‘exposure’ variables. However, unlike in randomized trials, which are often viewed as the gold standard to estimate controlled outcome differences (Zabor et al., 2020), certain observational scenarios preclude randomizing individuals to a particular experimental condition due to moral or biological restrictions. These include practical examples such as the timing of transplant waitlisting or the inability to ‘assign’ race to an individual (Peng et al., 2017; Li and Li, 2023). Our paper addresses this knowledge gap by proposing a strategy for incorporating survey weights from complex survey designs with propensity score methods to achieve covariate balance among heterogeneous groups, such as race.
Given the complex relationship between race, socioeconomic status (SES), and the porous collection of such data in the United States, the ability to make definitive statements on race differences in health outcomes is immensely challenging (Li and Li, 2023). Observational data can offer a wealth of information, however, many observational studies have their limitations in terms of confounding, covariate imbalance, and a lack of representation for generalizing results (Imai and Ratkovic, 2013; Pattanayak et al., 2011). Nationally representative surveys provide the opportunity to study individuals from a range of socioeconomic and cultural backgrounds using techniques such as stratified and clustered sampling methods, however researchers must consider the additional challenge of generalizing results outside the context of the sample. Such generalizations of group differences falter unless we adjust our comparison groups so that we have have similar levels of other covariates that are correlated with the comparison group and outcome of interest. This adjustment, which we term covariate balance can be accomplished a number of ways. In this paper, we use propensity scores. Rosenbaum and Rubin (1983) originally proposed propensity scores as a method of contrasting the conditional probability of experimental group assignment while adjusting for baseline covariates (Rosenbaum and Rubin, 1983). In a typical causal framework, the ability to reduce biases due to confounding by balancing covariate differences is a task that relies on several assumptions, namely (1) strong exchangeability of the treatment/exposure, (2) positivity of the treatment/exposure, and (3) stable unit treatment value. When transporting to a different sample or generalizing to a larger target population, we also require (4) strong selection exchangeability and (5) selection positivity. Rather than interpreting the difference in the observed outcomes as a causal effect, we seek to estimate a controlled descriptive comparison to adjust for differences in covariates between groups. Thus, we leverage propensity scores to compute what Li et al. (2013) call an average controlled difference (ACD). This quantity accounts for associations between variables of interest under weaker assumptions – positivity (1) and (5) and weak selection exchangeability (4) – but does not render a causal interpretation to the estimated group differences. While we later show that our proposed methods also target a causal quantity such as the population average treatment effect (PATE) under stronger assumptions, we do not require this for validity of our proposed estimation procedure, as one or more of these assumptions may be violated in observational data settings – particularly when measuring racial differences in health outcomes. We will later present an auxiliary example where we illustrate PATE estimation for an environmental exposure.
To further complicate things in the study of health disparities, it is often the case that the population characteristics of interest, such as race, are factors that influence the probability of being selected into the sample. For instance, the sampling mechanism for the National Health and Nutrition Examination Survey (NHANES) is partially determined by race and indicators of socioeconomic status. For example, NHANES oversamples non-Hispanic Black individuals and those individuals at or below 130% of the poverty level (Curtin et al., 2012). Previous studies have discussed using propensity score methods for analyzing complex survey data (Austin and Stuart, 2015; Lee et al., 2011; DuGoff et al., 2014; Fuentes et al., 2022), however, not when sample selection depends on race (or other variables commonly used for comparing outcome differences). For example, Rossen and Talih (2014) used propensity scores to assess disparities in weight among US children and adolescents from 2001 to 2010 with data from NHANES. However, there was not consensus on how to appropriately incorporate survey weights at the propensity score estimation stage.
There are several challenges in the problem we consider that are distinct from a traditional PATE estimation problem. A key assumption and function of the propensity score, traditionally, is conditional independence of the outcome and comparison group of interest conditional on the propensity score or any balancing score. Due to the complex nature of these relationships, it is not immediately clear which variables should be conditioned on in the propensity score versus the downstream outcome model. Second, it is often useful to express the research question and corresponding dependencies in terms of conceptual diagrams or directed acyclic graphs (such as Figure 1). In estimating the ACD, we include diagrams to demonstrate associations between the variables of interest. However, given the nature of the variables considered, we caution against interpreting these relationships as causal. If considering these diagrams broadly and even causally, in some instances it would be possible to open backdoor pathways when conditioning on selection into the survey, allowing for the possibility of violating assumptions that causal conclusions rely on. Addressing this in practice often relies on multi-stage modeling techniques such as inverse probability weighting (IPW) or direct standardization via ‘g-methods’ (see Richardson and Rotnitzky, 2014, for a more detailed history).
@@
In this paper, we introduce a strategy for incorporating sampling mechanisms, such as survey weights, and propensity scores to make statements on the controlled difference in outcomes when an individual’s probability of selection depends on the group membership variable of interest (herein, race; see Figure 1). We explore how the proposed methods compare to current analytic strategies. Our motivation comes from a research question addressing the relationship between race, SES, and telomere length, a biomarker for health measured in a particular nationally representative survey (NHANES). The remainder of this manuscript is structured as follows. Section 2 reviews related work in this area. Section 3 introduces and justifies our proposed methods for estimating the quantity of interest, which we will refer to as the ACD in the spirit of Li and Li (2023), and discusses differences with other common approaches. Section 4 provides simulation studies to support our recommendations. Section 5 introduces the data motivating this work and provides results for this example. We close with a summary discussion of the concepts outlined here and some final recommendations to researchers who may be working under similar considerations.
Figure 1: Conceptual diagrams. (A) Assumed variable relationships: $S$ = sample selection indicator; $A$ = group membership variable of interest; $X$ = covariates; $Y$ = observed outcome. (B) Conceptual diagram for our motivating data example: SES = Socioeconomic Status. (C) Directed acyclic graph for our auxiliary data example.
@@ 2 Related Work
To our knowledge, Zanutto (2006) and DuGoff et al. (2014) were among the first to discuss propensity score methods as an important tool for addressing confounding in observational studies that may include survey weights. Both authors suggest that, intuitively, since the purpose of incorporating propensity scores is to achieve in-sample balance rather than in the population, weighting the propensity score by the survey weight is unnecessary (Zanutto, 2006; DuGoff et al., 2014). The authors also recommend including survey weights as covariates in the propensity model, drawing a parallel between this intuition to the recommendation by Lumley (2011) of including these covariates to account for non-response in surveys (Lumley, 2011). However, this simple inclusion is often problematic in practice, as it may lead to separation in the data or large biases (Ridgeway et al., 2015). DuGoff et al. (2014) also addressed limitations regarding how to handle extreme weights, model misspecification, nonlinear propensity score estimation, and additional covariates in the models. Unfortunately, these recommendations are largely based on intuition, necessitating more rigorous arguments of consistency and conditional independence.
Subsequently, Ridgeway et al. (2015) began to formalize the discussion of propensity score analysis for complex survey designs. The authors demonstrated through proof and simulation that not appropriately accounting for the selection mechanism can potentially lead to biased estimates. A natural question is in what stages of the analysis (i.e., the propensity score and/or outcome model stage) should the survey weights be included? The authors recommended that researchers use survey weights in both stages of the based on their results. However, the simulation results pertaining to our scenario are inconclusive as to which approach achieves the lowest bias, and the derivations of consistency are not specific to this scenario. Further, they consider situations with only the population average treatment effect on the treated (PATT) as their target estimand (Ridgeway et al., 2015).
@@
Similarly, others have addressed the task of combining causal inference and survey methodology. Austin et al. (2018), for example, focused primarily on propensity score matching techniques in simulation. However, these results did not provide a clear answer with respect to which method of estimating the propensity score model, survey-weighted or not, was best. Dong et al. (2020) explored the performance of different propensity and outcome weighted models and suggested the use of survey weights at both stages, though they concluded the estimates from different strategies were not statistically different in their data example. Lenis et al. (2018) further studied the causal inference framework for estimating population versus sample effects and explored the impact of non-response mechanisms with matching techniques. Their simulation setting followed that of Austin et al. (2018) and conclude that the degree of misspecification impacted their analysis results and population balance. In our setting, the sample selection weights are supplied by NHANES, however the true sample selection mechanism is unknown. Further, there are many SES variables collected in the survey, which are potential covariates in selection and propensity score models. Here, considerations as to the limited degrees of freedom in the modeling process with the addition of survey weights, strata, and PSU’s may be especially important. Lastly, race is part of the selection mechanism in the NHANES design, which complicates the relationships underlying our motivating research question.
@@ 3.1 Notation
Let $Y$ denote our observed outcome, here telomere length. Our goal is to estimate the de-confounded, or controlled difference in mean telomere length between non-Hispanic White and non-Hispanic Black-identifying individuals in the United States. To that end, let $\mathcal{A}$ be the discrete set of potential comparison groups of interest, such that $a\in\mathcal{A}$ denotes a specific group. We are interested in comparing the values of $Y$ across levels of $A$. Our motivating analysis departs from the traditional causal literature in that $\mathcal{A}$ represents race, with $a=1$ denoting Black race and $a=0$ denoting White race, rather than a potential ‘treatment,’ per se (see Li and Li, 2023, for further discussion on using propensity scores for racial disparities analysis). Given the nature of our variables, and race in particular, this contrast of interest will primarily be viewed as a controlled comparison of the observed outcomes after balancing the two groups with respect to potential confounding variables (Li et al., 2013; Li and Li, 2023). As the lived experience of Black and White individuals differs significantly in terms of social determinants of health, we believe a controlled descriptive comparison can help inform the dynamics of these complex relationships (Li and Li, 2023). While race may not be interpreted as a modifiable treatment, we conceptualize race as an important construct worthy of study. Moreover, the methods described herein are equally applicable in settings where a participant’s sample selection is dependent on a treatment condition or exposure status in a more traditional causal inference framework (see Section 3.2 for more details).
Our observed data, $\mathcal{D}=\{(Y_{i},A_{i},X_{i},S_{i}=1);i=1,\ldots,n\}$, consist of $n$ independently and identically distributed copies of $(Y,A,X,S)$ sampled from a population distribution, $\mathcal{P}$, of $N$ individuals where $Y$ is the observed outcome, $A$ is a participant’s group membership, $X$ is a $p$-dimensional vector of covariates, and $S$ is an indicator for being selected into the sample. In our motivating example, $\Pr(S_{i}=1)$ is known for each individual $i$ in the super population through the explicit NHANES sampling design. There is also the possibility of having missing outcome values among sampled individuals, but this situation will not be explored here and is instead left to future work. Further, our assumed variable relationships are given in Figure 1. Here, we can see that our outcome ($Y$), the variable of interest ($A$), and the sample selection indicator ($S$) all depend on a particular set of covariates ($X$), and sample selection also depends on race.

### v2

@@ Abstract
In this work, we are motivated by the problem of estimating racial disparities in health outcomes, specifically the average controlled difference (ACD) in telomere length between Black and White individuals, using data from the National Health and Nutrition Examination Survey (NHANES). To do so, we build a propensity for race to properly adjust for other social determinants while characterizing the controlled effect of race on telomere length. Propensity score methods are broadly employed with observational data as a tool to achieve covariate balance, but how to implement them in complex surveys is less studied – in particular, when the survey weights depend on the group variable under comparison (as the NHANES sampling scheme depends on self-reported race). We propose identification formulas to properly estimate the ACD in outcomes between Black and White individuals, with appropriate weighting for both covariate imbalance across the two racial groups and generalizability. Via extensive simulation, we show that our proposed methods outperform traditional analytic approaches in terms of bias, mean squared error, and coverage when estimating the ACD for our setting of interest. In our data, we find that evidence of racial differences in telomere length between Black and White individuals attenuates after accounting for confounding by socioeconomic factors and utilizing appropriate propensity score and survey weighting techniques. Software to implement these methods and code to reproduce our results can be found in the R package svycdiff, available through the Comprehensive R Archive Network (CRAN) at cran.r-project.org/web/packages/svycdiff/, or in a development version on GitHub at github.com/salernos/svycdiff.
Keywords Complex Surveys  $\cdot$ Controlled Outcome Differences  $\cdot$ NHANES  $\cdot$ Propensity Scores  $\cdot$ Racial Disparities
@@ 1 Introduction
There is a deep and broad literature in the United States (US) documenting racial disparities along multiple socioeconomic axes, such as income and education (Needham et al., 2013; Rehkopf and Needham, 2019; Zota and VanNoy, 2021). These factors are interrelated, making it difficult to separate the effect of race from other sources. Further, these complex relationships between race and socioeconomic status (SES), and the porous collection of such data in the US, make it challenging to draw definitive conclusions about race differences in health outcomes (Li and Li, 2023). Observational studies can offer a wealth of information, but they are often limited by confounding, covariate imbalance, and a lack of representation (Imai and Ratkovic, 2013; Pattanayak et al., 2011). Nationally representative surveys, using techniques such as stratified and clustered sampling, allow us to study diverse populations, but generalizing results while accounting for confounding can be difficult due to their complex designs. Our motivating question focuses on the relationship between race, SES, and telomere length, a biomarker for health, using data from the National Health and Nutrition Examination Survey (NHANES): if we could hypothetically balance SES between Black and White individuals in a nationally representative sample, would we still see significant Black/White differences in telomere length?
This question is statistically complex because the population characteristic of interest (i.e., race) is correlated with SES, and both factors influence the probability of being selected into the sample. For NHANES, the sampling mechanism is partly determined by race and socioeconomic status, with an oversampling of non-Hispanic Black individuals and those at or below 130% of the poverty level (Curtin et al., 2012). To address this, we propose a strategy that appropriately accounts for both confounding and selection bias when estimating the controlled difference in outcomes between groups, where an individual’s probability of selection depends on their group membership and its confounders.
Rosenbaum and Rubin (1983) proposed using propensity scores to balance the distributions of covariates between comparison groups in the presence of confounding bias. In traditional causal inference, these comparison groups are analogous to randomized ‘treatments’ or ‘exposures,’ and using propensity scores to reduce confounding bias relies on several assumptions, namely exchangeability of the treatment/exposure, positivity of the treatment/exposure, and the stable unit treatment value assumption. However, certain observational scenarios preclude randomizing individuals due to moral or biological restrictions. These include practical examples such as the timing of transplant waitlisting or the inability to ‘assign’ race to an individual (Peng et al., 2017; Li and Li, 2023). Rather than interpreting differences in outcomes across these groups as causal effects, we seek a controlled descriptive comparison. Propensity scores can still be used to compute what Li et al. (2013) call an average controlled difference (ACD). This quantity achieves covariate balance under the weaker assumption of positivity, but it does not render a causal interpretation to the estimated group differences. We will refer to race as the ‘group’ (rather than ‘treatment’) of interest hereafter.
While Li and Li (2023) showed that propensity scores can be used in this way to study racial disparities, a gap remains in determining the ideal practice when the sampling design is complex. Many studies have discussed using propensity score methods to generalize inference from randomized trials (Stuart et al., 2011; Lim et al., 2018; Zeng et al., 2021), and complex surveys (Austin and Stuart, 2015; Lee et al., 2011; DuGoff et al., 2014; Fuentes et al., 2022), however, not when sample selection depends on the group of interest, such as race. For example, Rossen and Talih (2014) used propensity scores to assess disparities in weight among US children and adolescents from 2001 to 2010 with data from NHANES. However, there was no consensus on whether to incorporate survey weights when estimating the propensity score and/or outcome model. Our paper addresses this knowledge gap by proposing a strategy for incorporating survey weights from complex survey designs with propensity score methods to achieve covariate balance and generalizability, specifically when survey selection depends on the group of interest. We offer four approaches for estimating the ACD, which rely on multi-stage modeling techniques such as inverse probability weighting (IPW) or direct standardization via ‘g-methods’ (see Richardson and Rotnitzky, 2014) and only require the assumptions of positivity, selection positivity, and weak selection exchangeability. While we later show that our proposed methods also target a causal quantity such as the population average treatment effect (PATE) under stronger assumptions, we do not require this for validity of our proposed estimation procedure, as one or more of these assumptions may be violated in observational data settings – particularly when studying racial differences in health outcomes.
In this work, we introduce a novel framework that enhances existing approaches for estimating racial disparities in health outcomes in complex survey settings. Our contributions are threefold. First, we derive new identification formulas that provide theoretical insight into the estimation of controlled group differences under conditions of confounding and selection bias, addressing a critical gap in the literature where selection depends on the group of interest, such as race. Second, we present a comprehensive set of simulation settings that consider different data-generating mechanisms and allow for a thorough examination of the proposed methodology’s performance under various real-world scenarios. Lastly, we implement these methods within a comprehensive R package, enabling researchers to rigorously and efficiently apply these techniques to their own data. Together, these contributions offer a robust framework for analyzing health disparities using complex survey data.
@@
The rest of this manuscript is structured as follows. Section 2 reviews related work in this area. Section 3 introduces and justifies our proposed methods for estimating the quantity of interest, which we will refer to as the ACD in the spirit of Li and Li (2023), and discusses differences with other common approaches. Section 4 explores how the proposed methods compare to current analytic strategies and provides simulation studies to support our recommendations. Section 5 introduces the data motivating this work and provides results for this example. We then present an auxiliary example where we estimate the PATE for an environmental exposure. We close with a summary discussion of the concepts outlined here and some final recommendations to researchers who may be working under similar considerations.
@@ 2 Related Work
To our knowledge, Zanutto (2006) and DuGoff et al. (2014) were among the first to discuss using propensity score methods to address confounding in observational studies with survey weights. Both authors suggest that, intuitively, since the purpose of propensity scores weighting is to achieve in-sample balance, survey weighting the propensity score is unnecessary. The authors also recommend including survey weights as covariates in the propensity model, drawing a parallel between this intuition to the recommendation by Lumley (2011) of including these covariates to account for non-response in surveys. However, this simple inclusion is often problematic in practice, as it may lead to separation in the data or large biases (Ridgeway et al., 2015). DuGoff et al. (2014) also addressed limitations regarding how to handle extreme weights, model misspecification, nonlinear propensity score estimation, and additional model covariates. Such recommendations are largely based on intuition for our particular setup and require more rigorous study. Subsequently, Ridgeway et al. (2015) began to formalize the discussion of propensity score analysis for complex survey designs. The authors demonstrated through proof and simulation that not appropriately accounting for the selection mechanism can potentially lead to biased estimates. A natural question is in what stages of the analysis (i.e., the propensity score and/or outcome model stage) should the survey weights be included? The authors recommended that researchers use survey weights in both stages based on their results. However, the simulation results pertaining to our scenario are inconclusive as to which approach achieves the lowest bias, and the derivations of consistency are not specific to this scenario. Further, they consider only the population average treatment effect on the treated.
@@
Similarly, others have addressed the task of combining causal inference and survey methodology. Austin et al. (2018), for example, focused primarily on propensity score matching techniques in simulation. However, these results did not provide a clear answer with respect to which method of estimating the propensity score model, survey-weighted or not, was best. Dong et al. (2020) explored the performance of different propensity and outcome weighted models and suggested the use of survey weights at both stages, though they concluded the estimates from different strategies were not statistically different in their data example. Lenis et al. (2018) further studied the causal inference framework for estimating population versus sample effects and explored the impact of non-response mechanisms with matching techniques. Their simulation setting followed Austin et al. (2018) and concluded that the degree of misspecification impacted their analysis results and population balance. More recently, Yang et al. (2024) conducted simulations, also following the design of Austin et al. (2018), to compare various propensity score weighting approaches for estimating population-level treatment effects on a binary outcome with survey weighted data. The authors found that two-stage methods, with predicted outcomes weighted by survey weights, outperformed the other methods for estimating the population average treatment effect on the treated, but for population average treatment effects, the best performing method depended on the degree of model misspecification and propensity score overlap. McCaffrey et al. (2024) further extended the use of survey and attrition weights from binary treatment propensity score analyses to the continuous exposure setting. They first derive the conditions under which sampling or nonresponse weights must be incorporated into both the generalized propensity score (GPS) estimation and the subsequent outcome model to obtain unbiased estimates of a continuous dose-response curve. Through an extensive simulation study, they show that while including weights in both stages generally ensures against bias, in some settings (e.g., when the same covariates drive both treatment assignment and sample selection), unbiased estimation is possible without weighting the GPS itself. Conversely, naively omitting weights can lead to bias or inefficiency when selection and/or attrition depend on treatment or covariates. They further demonstrate that, although weighting both stages often reduces bias, it can inflate variance when weights are highly variable. In our setting, the sample selection weights are supplied by NHANES, and race is part of the selection mechanism in the NHANES design, which complicates the relationships underlying our motivating research question.
@@ 3.1 Notation
Let $Y$ denote our outcome, here telomere length. Our goal is to estimate the deconfounded, or controlled difference in mean telomere length between non-Hispanic White and non-Hispanic Black-identifying individuals in the US. To that end, let $\mathcal{A}$ be the discrete set of potential comparison groups of interest, such that $a\in\mathcal{A}$ denotes a specific group. We are interested in comparing the values of $Y$ across levels of $A$. Our motivating analysis departs from the traditional causal literature in that $\mathcal{A}$ represents race, with $a=1,0$ denoting Black versus White race, respectively, rather than a potential ‘treatment,’ per se. Given the nature of these variables, our contrast of interest will be viewed as a controlled comparison of the observed outcomes after balancing the two groups with respect to potential confounding variables (Li et al., 2013; Li and Li, 2023). As the lived experience of Black and White individuals differs significantly in terms of social determinants of health, we believe a controlled descriptive comparison can help inform the dynamics of these complex relationships (Li and Li, 2023). While race may not be interpreted as a modifiable treatment, the methods described herein are equally applicable in settings where sample selection depends on a treatment in a more traditional sense (see Section 3.2).
Our observed data are $\mathcal{O}=\{(Y_{i},A_{i},\boldsymbol{X}_{i});i=1,\ldots,n\}\overset{\text{i.i.d.}}{\sim}\mathcal{Q}:=\Pr(\ \cdot\mid S=1)$, where $\mathcal{Q}$ is the sample law for units selected into the analytic sample ($S_{i}=1$), $S\in\{0,1\}$ indicates selection, and $\boldsymbol{X}\in\mathbb{R}^{p}$ denotes our (potentially imbalanced) covariates. The underlying population law is $\mathcal{P}:=\Pr(\cdot)$ on $N$ individuals. In our motivating data, survey weights, and hence the design probabilities, $\Pr(S=1\mid\cdot)$, are known from the sampling plan. There is also the possibility of having missing outcome values among sampled individuals, but this situation will not be explored here and is instead left to future work. Our assumed variable relationships are given in Figure 1. Here our outcome ($Y$), the variable of interest ($A$), and the sample selection indicator ($S$) all depend on covariates ($\boldsymbol{X}$), and sample selection also depends on $A$. Unlike in a causal setting, where $\boldsymbol{X}$ consists of variables which affect both the treatment assignment and the outcome (for example, consider the DAG in the third panel of Figure 1), in the ACD setting, $\boldsymbol{X}$ consists of covariates which are unbalanced across $A$, and will bias the ACD estimate if not properly accounted for. Further, $\boldsymbol{X}$ depends on the research question. That is, $\boldsymbol{X}$ is the set of covariates for which we wish to balance to estimate the controlled difference between outcomes. To some extent, this answers the question of, if we balance the distribution of $\boldsymbol{X}$ between groups, what is the “residual difference” in outcomes? Here, if we do not account for the contribution of $\boldsymbol{X}$ to the group differences, then we cannot estimate a balanced, marginal effect of the comparison group membership.
Figure 1: Conceptual diagrams. (A) Assumed variable relationships: $S$ = sample selection indicator; AA = group membership variable of interest; $\boldsymbol{X}$ = covariates; $Y$ = observed outcome. (B) Conceptual diagram for our motivating data example: SES = Socioeconomic Status. (C) Directed acyclic graph for our auxiliary data example.